OpenStudy (anonymous):

Hi, I need help understanding how to solve this 3x3 System of Equations by elimination: 4x + 4y + z=24 2x - 4y + z = 0 5x - 4y - 5z = 12 I already know the values for each variable, I just want to know how to solve it myself

OpenStudy (anonymous):

Does it has to be by elimination?

OpenStudy (anonymous):

Yes please, the test will only have elimination problems and Cramer's rule. I already know how to use the latter

OpenStudy (anonymous):

Sorry, I don't like elimination :c

OpenStudy (anonymous):

All good

OpenStudy (anonymous):

4x+4y+z=24 (1) 2x-4y+z=0 (2) 5x-4y-5z= 12 (3) add (1) and (2) 6x+2z=24 3x+z=12 (4) add (1) and (3) 9x-4z=36 (5) multiply (4) by 4 and add in (5) 21x=48+36=84 x=84/21=4 put in (4) 3*4+z=12 z=12-12=0 put in (2) 2*4-4y+0=0 8=4y y=8/4=2

OpenStudy (anonymous):

In the calculator I used, I got y = -4 and z = 5

OpenStudy (anonymous):

let us check the solution put x=4,y=2 z=0 in any equation say (3) 5*4-4*2-5*0=12 20-8=12 or 12=12 which is true.

OpenStudy (anonymous):

Ah, cool. Thanks so much!

OpenStudy (anonymous):

yw